# Efficient and Transferable Agentic Knowledge Graph RAG via Reinforcement Learning

### v3

@@ 3.3 KG-R1 training with RL
Our goal is to find the KG-R1 agent’s policy $\pi_{\theta}(m_{t}\mid\text{context}_{t})$ that effectively retreive reasoning paths on $G$ via multi-turn interaction and generate the accurate answer $\hat{A}_{q}=A_{q}$. To this end, we optimize the base LLM with reinforcement learning, using a GRPO-style objective (DeepSeek-AI, 2025) in a multi-turn setting (Qian et al., 2025; Jin et al., 2025; Zeng et al., 2025). Our reward function combines action-level and outcome-based signals. The overview of the training procedure refers to Algorithm 1 (Appendix).
Reward Objective. To effectively train the KG-R1 agent, we combine verifiable turn rewards with outcome-level global rewards. Turn rewards ($r^{\text{turn}}_{t}$) provide local signals at each step as the sum of three components: (i) format validity $v_{\mathrm{fmt}}(m_{t})$ checks that $m_{t}$ contains both reasoning and a well-formed action $a_{t}\in\mathcal{A}$; (ii) KG query $v_{\mathrm{kg}}(a_{t},o_{t+1})$ checks that executing $a_{t}$ yields meaningful, schema-valid retrieval in $o_{t+1}$; and (iii) answer $v_{\mathrm{ans}}(m_{T})$ checks the final answer’s format/consistency on the final turn. the turn rewards are computed as follows with weight $w_{fmt},w_{kg},w_{ans}$:
@@
Global rewards summarize trajectory-level outcomes as the sum of the following: (i) a final-answer F1 score over the predicted set $\hat{A}_{q}$ vs. ground-truth answer set $A_{q}$; and (ii) a retrieval score $v_{ret}$ that is $1$ if any gold entity appears anywhere in the retrieved information along the executed reasoning path, and 0 otherwise.
$R^{\mathrm{global}}=w_{\mathrm{F1}}\cdot\mathrm{F1}\!\bigl(\hat{A}_{q},A_{q}\bigr)\;+\;w_{\mathrm{ret}}\cdot v_{ret}$ (2)
Group-relative turn-level credit assignment and optimization. To convert rewards into token-level credit, we use a group-relative, turn-level credit assignment inspired by Zeng et al. (2025). We collect $N$ rollouts per query $q$. For each episode $\tau^{(n)}$, we compute turn rewards $r^{\mathrm{turn},(n)}_{t}$ and a global reward $R^{\mathrm{global},(n)}$, and then form turn-specific returns with $\lambda$ as the normalization factor:
@@
$A_{t}^{(n)}=\frac{G_{t}^{(n)}-\bar{G}}{\sigma_{G}+\varepsilon}\quad\text{where}\quad\bar{G}=\tfrac{1}{|\mathcal{S}|}\!\sum_{(n,t)\in\mathcal{S}}G_{t}^{(n)},\ \sigma_{G}=\sqrt{\tfrac{1}{|\mathcal{S}|}\!\sum_{(n,t)\in\mathcal{S}}\!\big(G_{t}^{(n)}-\bar{G}\big)^{2}}.$
where $\varepsilon$ is a small constant for numerical stability.
RL update. Using turn-level credit $A_{t}^{(n)}$, we optimize the agent’s policy $\pi_{\theta}$ with a GRPO-style objective $\mathcal{J}$ :
$\mathcal{J}(\theta)=\mathbb{E}\!\Bigg[\sum_{n,t,i}\min\!\Big(\rho_{t,i}^{(n)}\,\widetilde{A}_{t,i}^{(n)},\ \operatorname{clip}\big(\rho_{t,i}^{(n)},\,1-\epsilon,\,1+\epsilon\big)\,\widetilde{A}_{t,i}^{(n)}\Big)\;-\;\beta\,\mathrm{KL}\!\big(\pi_{\theta}(\cdot\mid h_{t,i}^{(n)})\,\|\,\pi_{0}(\cdot\mid h_{t,i}^{(n)})\big)\Bigg]$
with $\widetilde{A}_{t,i}^{(n)}=m_{t,i}^{(n)}A_{t}^{(n)}$. $\pi_{\theta}$ is the current policy, and $\pi_{\theta_{\text{old}}}$ is the behavior policy used for sampling; $\pi_{0}$ is a fixed reference policy used only for KL regularization (weight $\beta$). $\rho_{t,i}^{(n)}=\exp\!\big(\log\pi_{\theta}(y_{t,i}^{(n)}\!\mid h_{t,i}^{(n)})-\log\pi_{\theta_{\text{old}}}(y_{t,i}^{(n)}\!\mid h_{t,i}^{(n)})\big)$ is the importance ratio that corrects off-policy sampling at token $i$. We maximize $\mathcal{J}(\theta)$ by gradient ascent. The proposed group-based credit assignment per turn produces stabilized and effective signals in the multi-turn interaction decisions. We provide comprehensive ablation studies of each component of our reward design in Section 4.3.
@@ 3.4 Cross-KG Transfer: Plug-and-Play
We propose a plug-and-play approach for KG-R1 that enables KG–RAG to operate across different knowledge graphs without retraining. Let $\pi_{\theta\mid\mathcal{D}}$ denote the policy of the KG-R1 agent trained on a KGQA dataset $\mathcal{D}=\{(q,G,A_{q})\}$. For a new KGQA dataset $\mathcal{D}^{*}=\{(q^{*},G^{*},A_{q^{*}})\}$, we replace the backend KG of the retrieval server with $G^{*}$ (plug) and evaluate $\pi_{\theta\mid\mathcal{D}}$ on $\mathcal{D}^{*}$ without any modification (play). This approach requires no task-specific fine-tuning, data relabeling, or architectural changes.
@@ 4 Experiments
We assess KG-R1 along two axes. (1) Performance and Efficiency on Trained KGs: We train KG-R1 on the training splits of KGQA benchmarks and measure answer quality and inference cost on their held-out test splits. (2) Cross-KG transferability: We test how well KG-R1 carries over to KGs out of training. In both parts, we compare against established baselines to support the KG-R1 results.
4.1 Performance and efficiency on Trained KGs
@@
Models. We use Qwen2.5-3B-it (Qwen et al., 2025) as our base model. For the other baseline methods, we use GPT-4o-mini, GPT-3.5, and Llama2-7B-it, following the setups used in prior work.
Datasets. We mainly evaluate KG-R1 in-domain on (1) WebQSP (Yih et al., 2016)—natural-language questions over Freebase, mostly 1–2 hop, using the official 3,098/1,639 train/test QA split; and (2) ComplexWebQuestions (CWQ) (Talmor & Berant, 2018)—compositional, multi-hop questions over Freebase with a 24,649/3,531 train/test QA split. For scalability, we extract 2-hop subgraphs for each question as in RoG (Luo et al., 2024). See Appendix B for detailed sources.
Metrics. Following prior work (Luo et al., 2024; Sun et al., 2024) to evaluate KGQA performance, we use F1 score to consider multiple valid predictions over answer sets and Hit@1 (1 if the gold answer appears in the single search, 0 otherwise). For efficiency, we report generated tokens (completion-only), measured end-to-end per query, aggregated across all turns ($1\ldots H$). This serves as a proxy for inference time and compute cost. For fair comparison, all token counts are computed with the Qwen2.5 tokenizer. We also report the number of modules to compare workflow complexity, and we analyze single-query latency and batched throughput on a single GPU node.
Baselines. We compare three classes of approaches: (1) LLM-only methods that do not access an external KG—Vanilla and Chain-of-Thought (CoT) prompting; (2) prior KG-augmented pipelines: RoG (Luo et al., 2024), which uses a fine-tuned modular workflow with LLaMA2-7B-it as the backbone, and ToG (Sun et al., 2024) and ReKnoS (Wang et al., 2025c), which use a prompt-based modular workflow with GPT-3.5; and (3) our method KG-R1. For LLM-only baselines, we employ LLM-as-Judge (Zheng et al., 2023) using gpt-4o-mini to semantically match predicted answers to ground truth. Full details for baselines are provided in Appendix C.
N-run beam search with KG-R1. Prior works often incorporated k-beam search (Lan & Jiang, 2020; Sun et al., 2024; Luo et al., 2024), retrieving k reasoning paths for enhanced search range on KGs and reporting improved performance. Following this idea, we propose KG-R1 N-run beam search : we run $N$ independent generation per question and collect predicted answer sets $\{\hat{A}_{q}^{(i)}\}_{i=1}^{N}$. We then take the union $\hat{A}_{q,\text{N-runs}}\;=\;\bigcup_{i=1}^{N}\hat{A}_{q}^{(i)},$ and compute F1 and Hit@1 on this union. This simple “$N$-beam” setup broadens search without extra orchestration, and cost scales roughly linearly with $N$ since runs can be executed in parallel.
Implementation details. We use VERL (Sheng et al., 2024) and vLLM-backed decoding (Kwon et al., 2023) for implementing the KG-R1 agent. Training uses distributed optimization with gradient checkpointing and offloading. Unless stated, we set $H=5$ turns and N=16N{=}16 rollouts per query. Additional implementation details (e.g., hyper-parameters) are provided in Appendix D.4.
4.1.1 KG-R1 Training Stability and Effectiveness
Figure 4 shows steady F1 improvement that plateaus after convergence during KG-R1 training on WebQSP and CWQ. The validation F1 scores track the training curve, indicating the agent learns generalization. Results are reproducible across three random seeds (Figure 6, Appendix), showing strong convergence with low variance. Overall, KG-R1 training yields stable, reproducible gains in end-to-end KG-RAG performance.
Figure 4: F1 score over KG-R1 training on WebQSP and CWQ for Qwen2.5-3B-it. Training (blue) and validation (red).
4.1.2 Inference Accuracy and Efficiency Results
Table 1: Performance and efficiency comparison of KGQA methods on WebQSP and CWQ datasets. The max turn number set to $H=5$. All scores are percentages.
Method Model Modules WebQSP CWQ Performance Efficiency Performance Efficiency F1 Hit@1 Total Gen F1 Hit@1 Total Gen LLM-only Methods Vanilla Qwen2.5-3B-it 1 29.4 46.6 95 30 16.6 21.1 104 42 Vanilla Qwen2.5-7B-it 1 33.2 50.9 95 67 20.7 25.7 104 92 Vanilla LLaMA2-7B-it 1 37.4 54.5 114 255 20.7 24.8 123 255 COT Qwen2.5-3B-it 1 30.6 47.6 131 192 17.3 21.4 140 216 COT Qwen2.5-7B-it 1 35.3 53.5 131 194 22.5 27.1 140 212 COT LLaMA2-7B-it 1 33.8 51.6 165 255 19.0 23.1 174 255 LLM+KG Methods RoG LLaMA2-7B-it 2 70.8 85.7 1.2K 295 56.2 62.6 1.1K 266 ToG GPT-3.5 4 72.8 76.2 3.5K 529 52.9 57.1 3.7K 520 ToG 2.0 GPT-3.5 5 74.5 77.8 39.K 605 65.8 68.9 3.8K 650 ReKnoS GPT-4o-mini 3 73.7 81.1 3.1K 617 64.7 66.8 4.1K 752 KG-R1 (Our Methods) KG-R1 (1 run) Qwen2.5-3B-it 1 77.5 84.7 3.2K 302 70.9 73.8 3.3K 377 KG-R1 (3 runs) Qwen2.5-3B-it 1 85.8 91.7 9.7K 906 81.0 83.9 10.0K 1.1K
A1. KG-R1 demonstrates strong accuracy on trained KGs. Table 1 reports accuracy and efficiency. On both WebQSP and CWQ, a single run (N=1N{=}1) of KG-R1 achieves competitive F1 and Hit@1 compared with non-KG and prior KG methods, including those using larger foundation models. With three runs with beam search, KG-R1 surpasses prior LLM+KG systems that also use $k$-beam search by a clear margin (WebQSP: F1 $77.5{\to}85.8$, Hit@1 $84.7{\to}91.7$; CWQ: F1 $70.9{\to}81.0$, Hit@1 $73.8{\to}83.9$). Overall, these results show that a lightweight KG-R1 agent delivers strong KG-RAG performance relative to previous baselines.
A2. KG-R1 achieves efficient inference. Table 1 reports efficiency (generation tokens and number of modules). For a single run (N=1N{=}1), KG-R1 produces $\sim$60 generation tokens per turn (300–370 tokens at H=5H{=}5), substantially below prior prompt-based LLM+KG systems (e.g., ToG $\sim$520 and ReKnoS $\sim$600–650) and comparable to the fine-tuned method RoG ($\sim$250–300). This small generation-token budget, coupled with a lightweight model, implies improved throughput and latency. Under N=1N{=}1, the total tokens are comparable to prior methods, indicating a similar memory (KV-cache) footprint. With three-run beam search (N=3N{=}3), token usage scales nearly linearly, exposing a tunable accuracy–efficiency trade-off via $N$.
A3. Latency and throughput analysis. On a single NVIDIA H100 GPU, we measure (i) single-query end-to-end latency and (ii) batched throughput. The single-query latency averages $6.4\pm 1.5$ s per question. Batched throughput reaches $3.7$ samples per second at batch size 64. The results suggest the feasibility of KG-R1’s single-node deployment. See Appendix E.2 for the full results.
4.2 Cross-KG Transfer via Plug-and-Play
In this subsection, we evaluate the plug-and-play approach (Sec. 3.4) for zero-shot generalization of our KG-R1 agent on knowledge graphs outside its training distribution.
@@ 4.2 Cross-KG Transfer via Plug-and-Play
Datasets We evaluate the plug-and-play approach on diverse KGQA benchmarks spanning three categories of KGs. (1) Freebase type: SimpleQA (Bordes et al., 2015) and GrailQA (Gu et al., 2021), which share the Freebase KG but differ in question complexity and reasoning path distributions. (2) Wikidata type: T-REx (Elsahar et al., 2018) and QALD-10en (Usbeck et al., 2023) for cross-schema generalization. (3) Temporal reasoning benchmark: MultiTQ (Chen et al., 2023), which requires reasoning over the ICEWS temporal KG (Boschee et al., 2015) at day, month, and year granularities. Appendix B provides a summary of dataset sources, schema, and evaluation splits.
Baselines For LLM-only methods, we evaluate (i) Vanilla and (ii) Chain-of-Thought (CoT) prompting with Qwen2.5–3B-it. For LLM+KG methods, we report results for RoG, ToG, and ReKnoS without modifying their modules (i.e prompts and models). In addition, for each KGQA benchmark, we include one or two recent state-of-the-art KG-specific methods (trained/tuned on the target dataset with task-specific pipelines) and use their published test-set scores. Full baseline details are provided in Appendix C.3.

### v5

@@ 3.3 KG-R1 Training with RL
Our goal is to learn a policy $\pi_{\theta}(m_{t}\mid\text{context}_{t})$ for the KG-R1 agent that interacts with the knowledge graph $G$ over multiple turns to incrementally retrieve informative reasoning paths and ultimately produce the correct answer $\hat{A}_{q}=A_{q}$. We optimize this policy using reinforcement learning with a GRPO-style objective [8], following recent advances in multi-turn agent optimization [19, 12, 32]. An overview of the training procedure is provided in Algorithm 1 in the Appendix˜E.
Reward Objective. To effectively train the KG-R1 agent, we combine verifiable turn rewards with outcome-level global rewards. Turn rewards ($r^{\text{turn}}_{t}$) provide local signals at each turn as the sum of three components: (i) format validity $v_{\mathrm{fmt}}(m_{t})$ checks that $m_{t}$ contains both reasoning and a well-formed action $a_{t}\in\mathcal{A}$; (ii) KG query $v_{\mathrm{kg}}(a_{t},o_{t+1})$ checks that executing $a_{t}$ yields meaningful, schema-valid retrieval in $o_{t+1}$; and (iii) answer consistency $v_{\mathrm{ans}}(m_{T})$ checks the final answer’s format/consistency on the final turn. The turn rewards are computed as follows with weights $w_{fmt}$, $w_{kg}$, and $w_{ans}$:
@@
Global rewards summarize trajectory-level outcomes as the sum of the following: (i) a binary answer correctness signal $\mathrm{H1}$ that is $1$ if the predicted answer $\hat{A}_{q}$ is included in the ground-truth answer set $A_{q}$, and 0 otherwise; and (ii) a retrieval score $v_{ret}$ that is $1$ if any ground-truth entity appears anywhere in the retrieved information along the executed reasoning path, and 0 otherwise.
$R^{\mathrm{global}}=w_{\mathrm{H1}}\cdot\mathrm{H1}\!\bigl(\hat{A}_{q},A_{q}\bigr)\;+\;w_{\mathrm{ret}}\cdot v_{ret}$ (2)
Group-relative Turn-level Credit Assignment and Optimization. To assign credit across multi-turn trajectories, we adopt a group-relative, turn-level credit assignment strategy [32]. We collect $N$ rollouts per query $q$. For each trajectory $\tau^{(n)}$, we compute turn rewards $r^{\mathrm{turn},(n)}_{t}$ and a global reward $R^{\mathrm{global},(n)}$, and combine them into a per-turn return with $\lambda$ as the normalization factor:
@@
$A_{t}^{(n)}=\frac{G_{t}^{(n)}-\bar{G}}{\sigma_{G}+\varepsilon}\quad\text{where}\quad\bar{G}=\tfrac{1}{|\mathcal{S}|}\!\sum_{(n,t)\in\mathcal{S}}G_{t}^{(n)},\ \sigma_{G}=\sqrt{\tfrac{1}{|\mathcal{S}|}\!\sum_{(n,t)\in\mathcal{S}}\!\big(G_{t}^{(n)}-\bar{G}\big)^{2}},$
where $\varepsilon$ is a constant for numerical stability. Intuitively, this measures each turn’s contribution relative to the group average, reinforcing above-average decisions and discouraging below-average ones.
RL update. Using the turn-level advantage $A_{t}^{(n)}$, we optimize the agent’s policy $\pi_{\theta}$ with a GRPO-style clipped objective $\mathcal{J}$:
$\mathcal{J}(\theta)=\mathbb{E}\!\Bigg[\sum_{n,t,i}\Big(\min\!\Big(\rho_{t,i}^{(n)}\widetilde{A}_{t,i}^{(n)},\ \operatorname{clip}\big(\rho_{t,i}^{(n)},1-\epsilon,1+\epsilon\big)\,\widetilde{A}_{t,i}^{(n)}\Big)-\beta\mathrm{KL}\!\big(\pi_{\theta}(\cdot\mid h_{t,i}^{(n)})\|\pi_{0}(\cdot\mid h_{t,i}^{(n)})\big)\Big)\Bigg],$
where $\widetilde{A}_{t,i}^{(n)}=m_{t,i}^{(n)}A_{t}^{(n)}$ applies the same turn-level advantage to all tokens within a turn via a mask $m_{t,i}^{(n)}$, $\operatorname{clip}$ bounds updates to improve stability, $n$, $t$, and $i$ index trajectories, turns, and tokens, respectively. The current policy is $\pi_{\theta}$, the behavior policy $\pi_{\theta_{\text{old}}}$ is used for sampling, and $\pi_{0}$ is a fixed reference policy used only for KL regularization (weight $\beta$). The importance ratio $\rho_{t,i}^{(n)}=\exp\!\big(\log\pi_{\theta}(y_{t,i}^{(n)}\!\mid h_{t,i}^{(n)})-\log\pi_{\theta_{\text{old}}}(y_{t,i}^{(n)}\!\mid h_{t,i}^{(n)})\big)$ corrects off-policy sampling at each token $i$. We maximize $\mathcal{J}(\theta)$ via gradient ascent. The proposed combination of structured rewards and group-relative credit assignment yields stable and informative learning signals for multi-turn decision-making, leading to improved reasoning quality and final answer correctness. We provide comprehensive ablation studies of each component of our reward design in Section 4.5.
@@ 3.4 Cross-KG Transfer: Plug-and-Play
We propose a plug-and-play approach for KG-R1 that enables KG-RAG to operate across different knowledge graphs without retraining. Let $\pi_{\theta\mid\mathcal{D}}$ denote the policy of the KG-R1 agent trained on a KGQA dataset $\mathcal{D}=\{(q,G,A_{q})\}$. For a new KGQA dataset $\mathcal{D}^{*}=\{(q^{*},G^{*},A_{q^{*}})\}$, we replace the backend KG of the retrieval server with $G^{*}$ (plug) and evaluate $\pi_{\theta\mid\mathcal{D}}$ on $\mathcal{D}^{*}$ without any modification (play). This approach requires no task-specific fine-tuning, data relabeling, or architectural changes.
@@ 4 Experiments
We evaluate KG-R1 through two research questions: RQ1 (Effectiveness and Efficiency): Can KG-R1 learn an effective and efficient KG-RAG policy when trained on a target KG?; RQ2 (Cross-KG Transferability): Does KG-R1 generalize to unseen KGs without additional training? For both research questions, we compare against established LLM-only and LLM+KG baselines to assess effectiveness, efficiency, and transferability.
4.1 Experimental Setup
@@
Models. We use Qwen2.5-3B-it [29] as our base model. For the other baseline methods, we use Qwen2.5-7B-it, Qwen2.5-72B-it, Qwen3-235B-A22B-Instruct-2507 [yang2025qwen3], and Llama2-7B-it (only for the fine-tuned planning in RoG [16]), following the setups used in prior work.
@@
Datasets. We mainly evaluate KG-R1 in-domain on (1) WebQSP [31]: natural-language questions over Freebase, mostly 1–2 hop questions, using the official 3,098/1,639 train/test QA split; and (2) ComplexWebQuestions (CWQ) [24]: compositional, multi-hop questions over Freebase with a 24,649/3,531 train/test QA split. For scalability, we extract 2-hop subgraphs for each question as in RoG [16]. See Appendix C for detailed sources. We evaluate the plug-and-play approach on diverse KGQA benchmarks spanning two categories of KGs: (1) Freebase type—SimpleQA [2] and GrailQA [11], which share the Freebase KG but differ in question complexity and reasoning path distributions; and (2) Wikidata type—T-REx [9] and QALD-10en [25] for cross-schema generalization. Appendix C provides a summary of dataset sources, schema, and evaluation splits.
Metrics. Following prior works [16, 23, 4, park2026prograg] to evaluate KGQA performance, we use Hits@1 (1 if the ground-truth answer appears in the single search, and 0 otherwise). For efficiency, we report total tokens and generated tokens (completion-only), measured end-to-end per query, aggregated across all turns. This serves as a proxy for inference time and compute cost. For fair comparison, all token counts are computed with the Qwen2.5 tokenizer. We also report the number of modules to compare workflow complexity, and we analyze single-query latency and batched throughput on a single NVIDIA H100 GPU.
Baselines. We compare three classes of approaches: (1) LLM-only methods that do not access an external KG: IO [brown2020gpt3] and Chain-of-Thought (CoT) [wei2022chain] prompting; (2) prior KG-augmented pipelines: RoG [16], which uses a fine-tuned modular workflow with LLaMA2-7B-it as the planner and we use the Qwen family as reasoner, and ToG [23] and ReKnoS [28], which use a prompt-based modular workflow; and (3) our method KG-R1. For LLM-only baselines, we employ LLM-as-Judge [34] using gpt-4o-mini [openai2024gpt4omini] to semantically match predicted answers to ground truth. We report the average and standard deviation over 5 runs for all models and baselines. Full details for the baselines are provided in Appendix D.
Implementation details. We use VERL [22] and vLLM-backed decoding [13] for implementing the KG-R1 agent. Training uses distributed optimization with gradient checkpointing and offloading. Unless stated otherwise, we set $H=5$ as the maximum number of turns and $N=16$ rollouts per query. Additional implementation details (e.g., hyperparameters) are provided in Appendix E.4.
Figure 4: Hits@1 over KG-R1 training on WebQSP and CWQ for Qwen2.5-3B-it.
4.2 Training Stability and Effectiveness
Figure 4 shows steady improvements in Hits@1 that gradually plateau as training converges for KG-R1 on both WebQSP and CWQ. The validation Hits@1 closely follows the training curve, indicating that the learned policy generalizes well to unseen data in the validation set without significant overfitting. Overall, KG-R1 training achieves stable and consistent gains in end-to-end KG-RAG performance.
4.3 RQ1: Effectiveness and Efficiency Results
Table 1: Performance and efficiency comparison of KGQA methods on WebQSP and CWQ datasets. The maximum turn number is set to $H=5$. All scores are percentages.
Datasets $\rightarrow$ WebQSP CWQ
Methods Modules Hits@1 (%) Total Tokens Gen Tokens Hits@1 (%) Total Tokens Gen Tokens
Qwen 2.5-7B IO 1 21.37 $\pm$ 0.07 151.59 $\pm$ 0.03 9.29 $\pm$ 0.03 16.63 $\pm$ 0.01 159.76 $\pm$ 0.01 8.57 $\pm$ 0.01
CoT 1 17.27 $\pm$ 0.09 260.43 $\pm$ 0.02 49.13 $\pm$ 0.02 14.95 $\pm$ 0.02 292.46 $\pm$ 0.01 72.27 $\pm$ 0.01
SC 1 19.26 $\pm$ 0.19 2098.01 $\pm$ 0.92 407.60 $\pm$ 0.92 17.63 $\pm$ 0.17 2364.62 $\pm$ 1.24 603.13 $\pm$ 1.24
RoG 2 79.89 $\pm$ 0.35 814.61 $\pm$ 0.70 246.95 $\pm$ 0.67 46.32 $\pm$ 0.49 1219.72 $\pm$ 0.61 259.46 $\pm$ 0.58
ToG 4 65.00 $\pm$ 4.63 6455.04 $\pm$ 753.60 941.52 $\pm$ 109.27 32.25 $\pm$ 0.43 8396.80 $\pm$ 87.06 1358.54 $\pm$ 17.24
PoG 5 64.81 $\pm$ 0.84 4474.30 $\pm$ 91.92 241.14 $\pm$ 5.17 32.00 $\pm$ 2.53 8293.67 $\pm$ 1454.76 369.84 $\pm$ 45.59
ReknoS 3 61.22 $\pm$ 5.46 2822.34 $\pm$ 209.87 558.43 $\pm$ 45.52 53.45 $\pm$ 3.03 2908.55 $\pm$ 222.47 602.08 $\pm$ 39.33
Qwen 2.5-72B IO 1 24.66 $\pm$ 0.07 153.50 $\pm$ 0.02 11.20 $\pm$ 0.02 22.20 $\pm$ 0.02 162.04 $\pm$ 0.03 10.85 $\pm$ 0.03
CoT 1 21.84 $\pm$ 0.04 271.46 $\pm$ 0.02 60.16 $\pm$ 0.02 22.92 $\pm$ 0.04 296.81 $\pm$ 0.04 76.62 $\pm$ 0.04
SC 1 22.62 $\pm$ 0.07 2177.75 $\pm$ 1.23 487.34 $\pm$ 1.23 24.76 $\pm$ 0.27 2385.73 $\pm$ 0.51 624.24 $\pm$ 0.51
RoG 2 81.63 $\pm$ 3.33 812.32 $\pm$ 7.81 244.53 $\pm$ 7.79 45.05 $\pm$ 2.87 1205.48 $\pm$ 2.51 245.34 $\pm$ 2.51
ToG 4 74.47 $\pm$ 0.25 6810.27 $\pm$ 39.73 942.91 $\pm$ 5.80 43.18 $\pm$ 1.02 9024.77 $\pm$ 271.68 1358.32 $\pm$ 41.02
PoG 5 72.88 $\pm$ 6.57 4759.10 $\pm$ 210.29 316.25 $\pm$ 16.85 43.62 $\pm$ 1.38 7969.21 $\pm$ 1631.58 503.29 $\pm$ 116.35
ReknoS 3 75.23 $\pm$ 9.25 2472.90 $\pm$ 115.94 488.44 $\pm$ 39.87 60.88 $\pm$ 2.54 2278.71 $\pm$ 196.08 548.52 $\pm$ 27.79
Qwen 3-235B IO 1 24.01 $\pm$ 0.18 136.05 $\pm$ 0.06 14.75 $\pm$ 0.06 20.22 $\pm$ 0.09 147.36 $\pm$ 0.16 17.17 $\pm$ 0.16
CoT 1 26.94 $\pm$ 0.22 274.66 $\pm$ 0.32 84.36 $\pm$ 0.32 24.38 $\pm$ 0.33 317.42 $\pm$ 0.31 118.23 $\pm$ 0.31
SC 1 27.60 $\pm$ 0.15 2200.20 $\pm$ 1.26 677.79 $\pm$ 1.26 26.13 $\pm$ 0.23 2551.71 $\pm$ 0.78 958.22 $\pm$ 0.78
RoG 2 81.68 $\pm$ 0.24 960.58 $\pm$ 5.90 254.31 $\pm$ 0.11 48.94 $\pm$ 0.39 1217.51 $\pm$ 10.26 252.45 $\pm$ 5.18
ToG 4 75.67 $\pm$ 0.37 6831.66 $\pm$ 28.39 996.12 $\pm$ 5.21 39.82 $\pm$ 0.73 5389.00 $\pm$ 33.01 802.97 $\pm$ 9.84
PoG 5 79.13 $\pm$ 1.89 6818.26 $\pm$ 908.76 418.26 $\pm$ 49.36 44.69 $\pm$ 0.27 10177.37 $\pm$ 1859.27 548.77 $\pm$ 82.56
ReknoS 3 84.73 $\pm$ 1.78 2795.53 $\pm$ 139.20 592.38 $\pm$ 4.18 65.63 $\pm$ 4.86 2677.92 $\pm$ 281.97 684.38 $\pm$ 14.06
KG-R1 (Ours) Qwen2.5-3B 1 82.75 $\pm$ 0.04 3237.17 $\pm$ 30.34 308.67 $\pm$ 0.56 65.31 $\pm$ 0.14 3205.95 $\pm$ 42.34 311.14 $\pm$ 0.52
KG-R1 demonstrates strong accuracy on trained KGs. Table 1 reports Hits@1 and token efficiency on WebQSP and CWQ. Despite being trained from a lightweight Qwen2.5-3B base model, KG-R1 achieves the strongest overall performance among all compared methods on WebQSP, reaching a Hits@1 of 82.75, outperforming RoG built on Qwen2.5-7B (79.89), Qwen2.5-72B (81.63), and even Qwen3-235B (81.68). It also exceeds or is comparable to stronger multi-module baselines such as ToG (75.67), PoG (79.13), and ReKnoS (84.73) while using a substantially smaller backbone. On the more compositional CWQ, KG-R1 achieves a Hits@1 of 65.31, improving over RoG by a large margin across all scales (46.32–48.94), and surpassing ReKnoS built on Qwen2.5-7B (53.45), Qwen2.5-72B (60.88), and Qwen3-235B (55.63). Notably, KG-R1 remains highly competitive with ReKnoS built on Qwen3-235B (65.31 vs. 65.63 on CWQ), despite using a Qwen2.5-3B backbone that is nearly two orders of magnitude smaller. This highlights that reinforcement learning over structured KG actions can substantially improve reasoning effectiveness beyond what is achievable by model scaling alone.
[…]
Single-agent design reduces system complexity. Beyond token efficiency, KG-R1 simplifies the KG-RAG pipeline into a single-agent architecture. Prior methods typically require 2–5 modules (e.g., retrieval, planning, verification, and answering), whereas KG-R1 performs retrieval and reasoning within a unified policy using only one module. This reduces deployment complexity and avoids error propagation across stages. Despite this simplification, KG-R1 remains competitive or superior to multi-module systems across both datasets, showing that structured RL can effectively consolidate KG interaction and reasoning into a lightweight unified agent.
Latency and throughput analysis. On a single NVIDIA H100 GPU, we measure (i) single-query end-to-end latency and (ii) batched throughput. The single-query latency averages $6.4\pm 1.5$ s per question, and the batched throughput reaches $3.7$ samples per second with a batch size of 64. The results suggest the feasibility of KG-R1’s single-node deployment. See Appendix F.1 for the full results.
4.4 RQ2: Cross-KG Transferability